# Fractal Patterns May Unravel the Intelligence in Next-Token Prediction

### v1

@@ 4 Related Works
The statistical attributes of human language have long piqued scholarly curiosity, such as One example is Zipf’s law, which Shannon leveraged to estimate the entropy of English to be around 1 bit per letter (Shannon, 1951), but his calculation did not consider second-order statistics. More recently,  Eftekhari (2006) proposed a refinement to Zipf’s law, suggesting its application to letters rather than words. Another related result is Heap’s law, which states that the number of unique words in a document is a power law function of the document’s length (Heaps, 1978). However, both Zipf’s and Heap’s laws are invariant to the semantic ordering of text, so they do not capture important aspects, such as long-range dependence (LRD) (Najafi & Darooneh, 2015).
@@
In terms of self-similarity in language, the Menzerath-Altmann law stipulates a self-similar behavior in the following sense: when the size of a language construct increases, the size of its constituents decreases, and this happens at all scales (Najafi & Darooneh, 2015; Andres, 2009). In Ausloos (2012), the authors model texts as a time series by replacing a word with its length. After that, they study the fractal behavior of language. However, replacing a word with its length is invalid because it is not translation-independent (i.e. one could map every word to an arbitrary token, including tokens of equal length). In our work, we model language as a time series of bits calculated from conditional entropies, reflecting the structure of the language itself.
@@
In Najafi & Darooneh (2015), the authors define a fractal dimension for each word. Informally, they examine the recurrence of a single, predetermined word in texts as an ON/OFF time series, similar to the approach used in Altmann et al. (2012). However, this is only applicable to individual words and cannot model higher-level clauses. For instance, it does not distinguish between the word “time” in the phrase “once upon a time” and the word “time” in “space and time.” Kokol & Podgorelec (2000) estimate LRD in natural language, and suggest that its LRD is close to that of pure noise! They conjecture this was due to the use of ASCII encoding. In computer languages, they observe LRD and suggest this is because computer languages are formal.
@@
Besides the above concerns in prior studies that examined the self-similar structure in language, another concern is that they sometimes give extremely large values of the fractal dimension, sometimes even exceeding 10 (Andres, 2009). Such values are difficult to interpret because classical definitions of the fractal dimension restrict its value to the range $[1,2]$ for time series. We do not observe such issues in our analysis. In our case, $\mathrm{D}=1.41\pm 0.08$.
@@ 5 Concluding Remarks
In this work, we highlight intriguing insights into the underlying fractal structure of language and how it may be interconnected with the intelligent behavior of LLMs. Our formalism quantifies properties of language that may have been suspected, but not previously formally shown. In particular, the need in LLMs to balance between short- and long-term contexts is reflected in the self-similar structure of language, while long-range dependence is quantifiable using the Hurst parameter. For instance, the absence of LRD in DM-Mathematics is reflected in its Hurst parameter of $\mathrm{H}\approx 0.5$. Interestingly, the estimated median Hurst value of $\mathrm{H}=0.70\pm 0.09$ in language reflects an intriguing balance between predictability and noise that is similar to many other phenomena, and combining both $\mathrm{H}$ with BPB together yields a stronger predictor of downstream performance. We carry out an extensive comparative analysis across different domains and model architectures, revealing that fractal parameters are generally robust. We hope that future research can further probe into these fractal properties, unearthing deeper understandings of the relation between intelligence and language.
6 Acknowledgement
@@
7 Potential Broader Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Experiment Details
All of our experiments are conducted in JAX/Flax (Bradbury et al., 2018) using the open source T5X framework (Roberts et al., 2022).
T5 baselines in Table 2 and 3 are pretrained from scratch using the open source T5.1.1 decoder-only architecture from the T5X library. [^1] https://github.com/google-research/t5x/tree/main/t5x/examples/decoder_only/models. We pretrain using a causal language modeling objective over the C4 corpus with the default T5 vocabulary as per Raffel et al. (2019). Training is done for 500k steps with a sequence length of 1024 and batch size of 512, resulting in a total of 262B tokens seen during pretraining. We optimize our model with the Adafactor (Shazeer & Stern, 2018) optimizer with an inverse square root learning rate schedule, 1k warmup steps, and an initial learning rate of 1e-2. Models are trained using 256 TPUv5e chips (Jouppi et al., 2020).
T5 context length ablation experiments in Table 4 are trained with the same pretraining objective but over the SlimPajama-627B corpus (Soboleva et al., 2023) and using a modified version of the T5 vocabulary that preserves whitespace and introduces byte-fallback for out of vocabulary tokens. This is similar to Chowdhery et al. (2022), but preserving the original T5 vocabulary. Models with sequence lengths 2048, 4096, 8192 are trained with batch sizes of 512, 256, and 128 respectively to preserve the number of tokens seen per batch and overall training steps. We train all models for 100k steps, using the same learning rate schedule described above. Hence, all models observe 100B tokens.
@@ Appendix B Full Results
In this section, we provide the full list of parameters calculated for each combination of LLM and domain. We use bootstrapping (Efron & Tibshirani, 1994) to estimate the error margin.
@@
Model
OpenWebText2
Github
FreeLaw
Pile-CC
Wikipedia
PubMed
Mathematics
ArXiv
T5-Decoder-110M
2.89
1.82
2.45
2.88
2.80
2.36
2.28
2.70
T5-Decoder-340M
2.60
1.56
2.14
2.62
2.52
2.08
2.10
2.42
T5-Decoder-1B
2.38
1.37
1.91
2.41
2.29
1.88
2.00
2.19
T5-Decoder-5B
2.19
1.22
1.73
2.25
2.11
1.73
1.91
2.01
PaLM1-8B
2.26
0.79
1.66
2.36
2.08
1.89
1.40
2.08
PaLM1-62B
2.02
0.62
1.44
2.14
1.80
1.68
1.30
1.83
PaLM1-540B
1.88
0.54
1.33
2.01
1.58
1.57
1.25
1.68
PaLM2-XXS
2.37
0.87
1.77
2.46
2.17
1.96
1.38
1.96
PaLM2-XS
2.12
0.73
1.53
2.22
1.92
1.72
1.27
1.72
PaLM2-S
1.95
0.60
1.37
2.06
1.71
1.57
1.19
1.55
PaLM2-M
1.88
0.56
1.31
1.99
1.59
1.51
1.12
1.48
PaLM2-L
1.75
0.46
1.23
1.88
1.22
1.43
1.08
1.36
@@
Model
OpenWebText2
Github
FreeLaw
Pile-CC
Wikipedia
PubMed
Mathematics
ArXiv
T5-Decoder-110M
$0.58\pm 0.04$
$0.67\pm 0.03$
$0.51\pm 0.02$
$0.54\pm 0.07$
$0.59\pm 0.04$
$0.59\pm 0.03$
$0.51\pm 0.04$
$0.58\pm 0.05$
T5-Decoder-340M
$0.52\pm 0.03$
$0.59\pm 0.05$
$0.63\pm 0.04$
$0.58\pm 0.04$
$0.61\pm 0.03$
$0.61\pm 0.03$
$0.48\pm 0.04$
$0.61\pm 0.05$
T5-Decoder-1B
$0.54\pm 0.01$
$0.66\pm 0.11$
$0.61\pm 0.06$
$0.57\pm 0.06$
$0.59\pm 0.05$
$0.60\pm 0.02$
$0.50\pm 0.03$
$0.63\pm 0.02$
T5-Decoder-5B
$0.51\pm 0.04$
$0.70\pm 0.04$
$0.60\pm 0.04$
$0.58\pm 0.02$
$0.58\pm 0.03$
$0.57\pm 0.02$
$0.45\pm 0.02$
$0.67\pm 0.05$
PaLM1-8B
$0.56\pm 0.03$
$0.67\pm 0.05$
$0.63\pm 0.05$
$0.58\pm 0.01$
$0.55\pm 0.04$
$0.62\pm 0.03$
$0.50\pm 0.03$
$0.68\pm 0.07$
PaLM1-62B
$0.49\pm 0.03$
$0.65\pm 0.09$
$0.63\pm 0.09$
$0.57\pm 0.03$
$0.63\pm 0.05$
$0.61\pm 0.04$
$0.48\pm 0.05$
$0.68\pm 0.03$
PaLM1-540B
$0.51\pm 0.04$
$0.68\pm 0.09$
$0.64\pm 0.05$
$0.58\pm 0.04$
$0.67\pm 0.03$
$0.64\pm 0.08$
$0.48\pm 0.03$
$0.65\pm 0.04$
PaLM2-XXS
$0.53\pm 0.02$
$0.61\pm 0.05$
$0.58\pm 0.04$
$0.60\pm 0.04$
$0.57\pm 0.05$
$0.61\pm 0.03$
$0.52\pm 0.02$
$0.70\pm 0.04$
PaLM2-XS
$0.54\pm 0.04$
$0.57\pm 0.06$
$0.58\pm 0.03$
$0.56\pm 0.04$
$0.60\pm 0.04$
$0.57\pm 0.06$
$0.45\pm 0.02$
$0.73\pm 0.06$
PaLM2-S
$0.55\pm 0.02$
$0.55\pm 0.15$
$0.59\pm 0.02$
$0.54\pm 0.08$
$0.65\pm 0.04$
$0.58\pm 0.05$
$0.49\pm 0.04$
$0.61\pm 0.03$
PaLM2-M
$0.58\pm 0.02$
$0.62\pm 0.06$
$0.59\pm 0.04$
$0.60\pm 0.05$
$0.70\pm 0.03$
$0.56\pm 0.04$
$0.46\pm 0.04$
$0.62\pm 0.05$
PaLM2-L
$0.53\pm 0.05$
$0.60\pm 0.05$
$0.61\pm 0.05$
$0.56\pm 0.03$
$0.62\pm 0.02$
$0.60\pm 0.07$
$0.42\pm 0.03$
$0.70\pm 0.03$
@@
Model
OpenWebText2
Github
FreeLaw
Pile-CC
Wikipedia
PubMed
Mathematics
ArXiv
T5-Decoder-110M
$0.63\pm 0.00$
$0.82\pm 0.01$
$0.62\pm 0.01$
$0.67\pm 0.01$
$0.62\pm 0.01$
$0.65\pm 0.00$
$0.54\pm 0.01$
$0.68\pm 0.01$
T5-Decoder-340M
$0.63\pm 0.01$
$0.82\pm 0.01$
$0.62\pm 0.00$
$0.67\pm 0.00$
$0.62\pm 0.01$
$0.64\pm 0.01$
$0.54\pm 0.00$
$0.67\pm 0.01$
T5-Decoder-1B
$0.63\pm 0.01$
$0.83\pm 0.01$
$0.63\pm 0.01$
$0.67\pm 0.00$
$0.62\pm 0.01$
$0.64\pm 0.00$
$0.54\pm 0.00$
$0.67\pm 0.00$
T5-Decoder-5B
$0.63\pm 0.01$
$0.82\pm 0.00$
$0.62\pm 0.01$
$0.67\pm 0.01$
$0.62\pm 0.01$
$0.64\pm 0.01$
$0.54\pm 0.00$
$0.67\pm 0.00$
PaLM1-8B
$0.65\pm 0.01$
$0.81\pm 0.01$
$0.66\pm 0.00$
$0.68\pm 0.01$
$0.66\pm 0.00$
$0.65\pm 0.01$
$0.57\pm 0.00$
$0.69\pm 0.01$
PaLM1-62B
$0.66\pm 0.01$
$0.80\pm 0.00$
$0.67\pm 0.01$
$0.69\pm 0.01$
$0.68\pm 0.00$
$0.65\pm 0.00$
$0.57\pm 0.00$
$0.70\pm 0.00$
PaLM1-540B
$0.67\pm 0.00$
$0.79\pm 0.01$
$0.68\pm 0.00$
$0.69\pm 0.01$
$0.71\pm 0.01$
$0.65\pm 0.01$
$0.56\pm 0.00$
$0.70\pm 0.01$
PaLM2-XXS
$0.65\pm 0.01$
$0.81\pm 0.01$
$0.65\pm 0.01$
$0.68\pm 0.01$
$0.66\pm 0.01$
$0.65\pm 0.01$
$0.58\pm 0.00$
$0.71\pm 0.01$
PaLM2-XS
$0.65\pm 0.01$
$0.81\pm 0.01$
$0.66\pm 0.01$
$0.68\pm 0.01$
$0.67\pm 0.00$
$0.65\pm 0.00$
$0.56\pm 0.01$
$0.71\pm 0.01$
PaLM2-S
$0.67\pm 0.01$
$0.80\pm 0.01$
$0.66\pm 0.01$
$0.69\pm 0.00$
$0.68\pm 0.01$
$0.65\pm 0.01$
$0.54\pm 0.00$
$0.71\pm 0.00$
PaLM2-M
$0.67\pm 0.01$
$0.80\pm 0.01$
$0.67\pm 0.01$
$0.70\pm 0.01$
$0.70\pm 0.01$
$0.65\pm 0.01$
$0.52\pm 0.01$
$0.72\pm 0.01$
PaLM2-L
$0.68\pm 0.01$
$0.79\pm 0.01$
$0.68\pm 0.00$
$0.70\pm 0.00$
$0.74\pm 0.01$
$0.65\pm 0.00$
$0.50\pm 0.01$
$0.72\pm 0.01$
@@
Model
OpenWebText2
Github
FreeLaw
Pile-CC
Wikipedia
PubMed
Mathematics
ArXiv
T5-Decoder-110M
$0.44\pm 0.01$
$0.53\pm 0.00$
$0.42\pm 0.00$
$0.49\pm 0.01$
$0.45\pm 0.00$
$0.43\pm 0.00$
$0.33\pm 0.00$
$0.45\pm 0.00$
T5-Decoder-340M
$0.44\pm 0.02$
$0.53\pm 0.00$
$0.43\pm 0.00$
$0.49\pm 0.00$
$0.45\pm 0.01$
$0.43\pm 0.00$
$0.33\pm 0.00$
$0.45\pm 0.00$
T5-Decoder-1B
$0.43\pm 0.01$
$0.53\pm 0.00$
$0.43\pm 0.01$
$0.49\pm 0.01$
$0.45\pm 0.01$
$0.42\pm 0.00$
$0.33\pm 0.00$
$0.45\pm 0.01$
T5-Decoder-5B
$0.43\pm 0.01$
$0.53\pm 0.00$
$0.44\pm 0.00$
$0.49\pm 0.01$
$0.45\pm 0.00$
$0.42\pm 0.00$
$0.34\pm 0.00$
$0.45\pm 0.00$
PaLM1-8B
$0.45\pm 0.00$
$0.51\pm 0.00$
$0.46\pm 0.00$
$0.49\pm 0.01$
$0.48\pm 0.01$
$0.44\pm 0.01$
$0.34\pm 0.00$
$0.48\pm 0.01$
PaLM1-62B
$0.45\pm 0.00$
$0.50\pm 0.01$
$0.47\pm 0.00$
$0.49\pm 0.01$
$0.49\pm 0.00$
$0.44\pm 0.00$
$0.33\pm 0.00$
$0.48\pm 0.01$
PaLM1-540B
$0.46\pm 0.01$
$0.49\pm 0.01$
$0.47\pm 0.00$
$0.50\pm 0.01$
$0.50\pm 0.00$
$0.44\pm 0.00$
$0.33\pm 0.01$
$0.48\pm 0.00$
PaLM2-XXS
$0.44\pm 0.01$
$0.50\pm 0.00$
$0.45\pm 0.00$
$0.50\pm 0.01$
$0.48\pm 0.00$
$0.45\pm 0.00$
$0.34\pm 0.00$
$0.49\pm 0.00$
PaLM2-XS
$0.45\pm 0.01$
$0.50\pm 0.01$
$0.46\pm 0.01$
$0.49\pm 0.00$
$0.48\pm 0.00$
$0.44\pm 0.00$
$0.33\pm 0.01$
$0.49\pm 0.00$
PaLM2-S
$0.45\pm 0.00$
$0.49\pm 0.00$
$0.47\pm 0.00$
$0.50\pm 0.01$
$0.50\pm 0.01$
$0.44\pm 0.00$
$0.31\pm 0.00$
$0.49\pm 0.00$
PaLM2-M
$0.45\pm 0.01$
$0.49\pm 0.01$
$0.48\pm 0.01$
$0.50\pm 0.01$
$0.50\pm 0.00$
$0.44\pm 0.00$
$0.29\pm 0.00$
$0.49\pm 0.01$
PaLM2-L
$0.46\pm 0.01$
$0.49\pm 0.00$
$0.49\pm 0.00$
$0.50\pm 0.00$
$0.52\pm 0.00$
$0.44\pm 0.00$
$0.28\pm 0.00$
$0.49\pm 0.00$

### v2

@@ 4 Related Works and Directions for Future Research
The statistical attributes of human language have long piqued scholarly curiosity, such as One example is Zipf’s law, which Shannon leveraged to estimate the entropy of English to be around 1 bit per letter [59], but his calculation did not consider second-order statistics. More recently,  Eftekhari, [18] proposed a refinement to Zipf’s law, suggesting its application to letters rather than words. Another related result is Heap’s law, which states that the number of unique words is a power law function of the document’s length [29]. However, both Zipf’s and Heap’s laws are invariant to the semantic ordering of text, so they do not capture important aspects, such as long-range dependence (LRD) [48].
@@
In terms of self-similarity in language, the Menzerath-Altmann law stipulates a self-similar behavior in the following sense: when the size of a language construct increases, the size of its constituents decreases, and this happens at all scales [48, 4]. In Ausloos, [9], the authors model texts as a time series by replacing a word with its length. After that, they study the fractal behavior of language. However, as mentioned in [22], replacing a word with its length is invalid because it is not translation-independent (i.e. one could map every word to an arbitrary token, including tokens of equal length). In our work, we model language as a series of bits calculated from conditional entropies, reflecting the intrinsic structure of the language itself, inspired by findings in linguistics such as [28, 22, 40].
@@
In Najafi and Darooneh, [48], the authors define a fractal dimension for each word. Informally, they examine the recurrence of a single, predetermined word as a binary series, similar to the approach used in Altmann et al., [3]. However, this only applies to individual words and cannot model higher-level clauses. For instance, it does not distinguish between “time” in the phrase “once upon a time” and “time” in “space and time.” Kokol and Podgorelec, [35] estimate LRD in natural language, and suggest that its LRD is close to that of pure noise! They conjecture this was due to their use of ASCII encoding. In computer languages, they observe LRD and suggest it is because they are formal.
@@
Besides the above concerns in prior studies that examined the self-similar structure in language, another concern is that they sometimes give extremely large values of the fractal dimension, sometimes exceeding 10 [4]! Such values are difficult to interpret because the fractal dimension $\mathrm{D}$ should fall in $\mathrm{D}\in[1,2]$ for time series. We do not observe such issues in our analysis. In our case, $\mathrm{D}=1.41\pm 0.08$.
Limitations and Future Research.
Our analysis is currently limited to the English language so it may not apply to other languages that differ significantly. For instance, some languages such as Pirahã (spoken in the Amazon) do not have a recursive structure like most languages do [20]. We also do not model the semantic or lexical form of language. While our information-theoretic approach is well-founded and captures the intrinsic complexity of language, it does not account for the semantic nuances that contribute to meaning. Thirdly, self-similarity may explain why parameter sharing, such as in ALBERT [37], can be successful but exploiting self-similarity more directly in LLMs could lead to further optimizations. Exploring these aspects are promising directions for future research.
@@ 5 Concluding Remarks
In this work, we highlight intriguing insights into the underlying fractal structure of language and how it may be interconnected with the remarkable capabilities of LLMs. Our formalism quantifies properties of language that may have been suspected, but not previously formally shown. In particular, the need in LLMs to balance between short- and long-term contexts is reflected in the self-similar structure of language, while long-range dependence is quantifiable using the Hurst parameter. For instance, the absence of LRD in DM-Mathematics is reflected in its Hurst parameter of $\mathrm{H}\approx 0.5$. Interestingly, the estimated median Hurst value of $\mathrm{H}=0.70\pm 0.09$ in language reflects an intriguing balance between predictability and noise that is similar to many other phenomena, and combining both $\mathrm{H}$ with BPB together yields a stronger predictor of downstream performance. We carry out an extensive comparative analysis across different domains and model architectures, revealing that fractal parameters are generally robust. We hope that future research can further probe into these fractal properties, unearthing deeper understandings of the relation between intelligence and language.
Acknowledgement
@@ Appendix A Experiment Details
All of our experiments are conducted in JAX/Flax [10] using the open source T5X framework [55].
T5 baselines in Table 2 and 3 are pretrained from scratch using the open source T5.1.1 decoder-only architecture from the T5X library. [^2] https://github.com/google-research/t5x/tree/main/t5x/examples/decoder_only/models. We pretrain using a causal language modeling objective over the C4 corpus with the default T5 vocabulary as per Raffel et al., [54]. Training is done for 500k steps with a sequence length of 1024 and batch size of 512, resulting in a total of 262B tokens seen during pretraining. We optimize our model with the Adafactor [60] optimizer with an inverse square root learning rate schedule, 1k warmup steps, and an initial learning rate of 1e-2. Models are trained using 256 TPUv5e chips [32].
T5 context length ablation experiments in Table 3 are trained with the same pretraining objective but over the SlimPajama-627B corpus [61] and using a modified version of the T5 vocabulary that preserves whitespace and introduces byte-fallback for out of vocabulary tokens. This is similar to Chowdhery et al., [12], but preserving the original T5 vocabulary. Models with sequence lengths 2048, 4096, 8192 are trained with batch sizes of 512, 256, and 128 respectively to preserve the number of tokens seen per batch and overall training steps. We train all models for 100k steps, using the same learning rate schedule described above. Hence, all models observe 100B tokens.
@@ Appendix B Full Results
In this section, we provide the full list of parameters calculated for each combination of LLM and domain. We use bootstrapping [17] to estimate the error margin.
@@
Model OpenWebText2 Github FreeLaw Pile-CC Wikipedia PubMed Mathematics ArXiv
T5-Decoder-110M 2.89 1.82 2.45 2.88 2.80 2.36 2.28 2.70
T5-Decoder-340M 2.60 1.56 2.14 2.62 2.52 2.08 2.10 2.42
T5-Decoder-1B 2.38 1.37 1.91 2.41 2.29 1.88 2.00 2.19
T5-Decoder-5B 2.19 1.22 1.73 2.25 2.11 1.73 1.91 2.01
PaLM1-8B 2.26 0.79 1.66 2.36 2.08 1.89 1.40 2.08
PaLM1-62B 2.02 0.62 1.44 2.14 1.80 1.68 1.30 1.83
PaLM1-540B 1.88 0.54 1.33 2.01 1.58 1.57 1.25 1.68
PaLM2-XXS 2.37 0.87 1.77 2.46 2.17 1.96 1.38 1.96
PaLM2-XS 2.12 0.73 1.53 2.22 1.92 1.72 1.27 1.72
PaLM2-S 1.95 0.60 1.37 2.06 1.71 1.57 1.19 1.55
PaLM2-M 1.88 0.56 1.31 1.99 1.59 1.51 1.12 1.48
PaLM2-L 1.75 0.46 1.23 1.88 1.22 1.43 1.08 1.36
@@
Model OpenWebText2 Github FreeLaw Pile-CC Wikipedia PubMed Mathematics ArXiv
T5-Decoder-110M $0.58\pm 0.04$ $0.67\pm 0.03$ $0.51\pm 0.02$ $0.54\pm 0.07$ $0.59\pm 0.04$ $0.59\pm 0.03$ $0.51\pm 0.04$ $0.58\pm 0.05$
T5-Decoder-340M $0.52\pm 0.03$ $0.59\pm 0.05$ $0.63\pm 0.04$ $0.58\pm 0.04$ $0.61\pm 0.03$ $0.61\pm 0.03$ $0.48\pm 0.04$ $0.61\pm 0.05$
T5-Decoder-1B $0.54\pm 0.01$ $0.66\pm 0.11$ $0.61\pm 0.06$ $0.57\pm 0.06$ $0.59\pm 0.05$ $0.60\pm 0.02$ $0.50\pm 0.03$ $0.63\pm 0.02$
T5-Decoder-5B $0.51\pm 0.04$ $0.70\pm 0.04$ $0.60\pm 0.04$ $0.58\pm 0.02$ $0.58\pm 0.03$ $0.57\pm 0.02$ $0.45\pm 0.02$ $0.67\pm 0.05$
PaLM1-8B $0.56\pm 0.03$ $0.67\pm 0.05$ $0.63\pm 0.05$ $0.58\pm 0.01$ $0.55\pm 0.04$ $0.62\pm 0.03$ $0.50\pm 0.03$ $0.68\pm 0.07$
PaLM1-62B $0.49\pm 0.03$ $0.65\pm 0.09$ $0.63\pm 0.09$ $0.57\pm 0.03$ $0.63\pm 0.05$ $0.61\pm 0.04$ $0.48\pm 0.05$ $0.68\pm 0.03$
PaLM1-540B $0.51\pm 0.04$ $0.68\pm 0.09$ $0.64\pm 0.05$ $0.58\pm 0.04$ $0.67\pm 0.03$ $0.64\pm 0.08$ $0.48\pm 0.03$ $0.65\pm 0.04$
PaLM2-XXS $0.53\pm 0.02$ $0.61\pm 0.05$ $0.58\pm 0.04$ $0.60\pm 0.04$ $0.57\pm 0.05$ $0.61\pm 0.03$ $0.52\pm 0.02$ $0.70\pm 0.04$
PaLM2-XS $0.54\pm 0.04$ $0.57\pm 0.06$ $0.58\pm 0.03$ $0.56\pm 0.04$ $0.60\pm 0.04$ $0.57\pm 0.06$ $0.45\pm 0.02$ $0.73\pm 0.06$
PaLM2-S $0.55\pm 0.02$ $0.55\pm 0.15$ $0.59\pm 0.02$ $0.54\pm 0.08$ $0.65\pm 0.04$ $0.58\pm 0.05$ $0.49\pm 0.04$ $0.61\pm 0.03$
PaLM2-M $0.58\pm 0.02$ $0.62\pm 0.06$ $0.59\pm 0.04$ $0.60\pm 0.05$ $0.70\pm 0.03$ $0.56\pm 0.04$ $0.46\pm 0.04$ $0.62\pm 0.05$
PaLM2-L $0.53\pm 0.05$ $0.60\pm 0.05$ $0.61\pm 0.05$ $0.56\pm 0.03$ $0.62\pm 0.02$ $0.60\pm 0.07$ $0.42\pm 0.03$ $0.70\pm 0.03$
@@
Model OpenWebText2 Github FreeLaw Pile-CC Wikipedia PubMed Mathematics ArXiv
T5-Decoder-110M $0.63\pm 0.00$ $0.82\pm 0.01$ $0.62\pm 0.01$ $0.67\pm 0.01$ $0.62\pm 0.01$ $0.65\pm 0.00$ $0.54\pm 0.01$ $0.68\pm 0.01$
T5-Decoder-340M $0.63\pm 0.01$ $0.82\pm 0.01$ $0.62\pm 0.00$ $0.67\pm 0.00$ $0.62\pm 0.01$ $0.64\pm 0.01$ $0.54\pm 0.00$ $0.67\pm 0.01$
T5-Decoder-1B $0.63\pm 0.01$ $0.83\pm 0.01$ $0.63\pm 0.01$ $0.67\pm 0.00$ $0.62\pm 0.01$ $0.64\pm 0.00$ $0.54\pm 0.00$ $0.67\pm 0.00$
T5-Decoder-5B $0.63\pm 0.01$ $0.82\pm 0.00$ $0.62\pm 0.01$ $0.67\pm 0.01$ $0.62\pm 0.01$ $0.64\pm 0.01$ $0.54\pm 0.00$ $0.67\pm 0.00$
PaLM1-8B $0.65\pm 0.01$ $0.81\pm 0.01$ $0.66\pm 0.00$ $0.68\pm 0.01$ $0.66\pm 0.00$ $0.65\pm 0.01$ $0.57\pm 0.00$ $0.69\pm 0.01$
PaLM1-62B $0.66\pm 0.01$ $0.80\pm 0.00$ $0.67\pm 0.01$ $0.69\pm 0.01$ $0.68\pm 0.00$ $0.65\pm 0.00$ $0.57\pm 0.00$ $0.70\pm 0.00$
PaLM1-540B $0.67\pm 0.00$ $0.79\pm 0.01$ $0.68\pm 0.00$ $0.69\pm 0.01$ $0.71\pm 0.01$ $0.65\pm 0.01$ $0.56\pm 0.00$ $0.70\pm 0.01$
PaLM2-XXS $0.65\pm 0.01$ $0.81\pm 0.01$ $0.65\pm 0.01$ $0.68\pm 0.01$ $0.66\pm 0.01$ $0.65\pm 0.01$ $0.58\pm 0.00$ $0.71\pm 0.01$
PaLM2-XS $0.65\pm 0.01$ $0.81\pm 0.01$ $0.66\pm 0.01$ $0.68\pm 0.01$ $0.67\pm 0.00$ $0.65\pm 0.00$ $0.56\pm 0.01$ $0.71\pm 0.01$
PaLM2-S $0.67\pm 0.01$ $0.80\pm 0.01$ $0.66\pm 0.01$ $0.69\pm 0.00$ $0.68\pm 0.01$ $0.65\pm 0.01$ $0.54\pm 0.00$ $0.71\pm 0.00$
PaLM2-M $0.67\pm 0.01$ $0.80\pm 0.01$ $0.67\pm 0.01$ $0.70\pm 0.01$ $0.70\pm 0.01$ $0.65\pm 0.01$ $0.52\pm 0.01$ $0.72\pm 0.01$
PaLM2-L $0.68\pm 0.01$ $0.79\pm 0.01$ $0.68\pm 0.00$ $0.70\pm 0.00$ $0.74\pm 0.01$ $0.65\pm 0.00$ $0.50\pm 0.01$ $0.72\pm 0.01$
@@
Model OpenWebText2 Github FreeLaw Pile-CC Wikipedia PubMed Mathematics ArXiv
T5-Decoder-110M $0.44\pm 0.01$ $0.53\pm 0.00$ $0.42\pm 0.00$ $0.49\pm 0.01$ $0.45\pm 0.00$ $0.43\pm 0.00$ $0.33\pm 0.00$ $0.45\pm 0.00$
T5-Decoder-340M $0.44\pm 0.02$ $0.53\pm 0.00$ $0.43\pm 0.00$ $0.49\pm 0.00$ $0.45\pm 0.01$ $0.43\pm 0.00$ $0.33\pm 0.00$ $0.45\pm 0.00$
T5-Decoder-1B $0.43\pm 0.01$ $0.53\pm 0.00$ $0.43\pm 0.01$ $0.49\pm 0.01$ $0.45\pm 0.01$ $0.42\pm 0.00$ $0.33\pm 0.00$ $0.45\pm 0.01$
T5-Decoder-5B $0.43\pm 0.01$ $0.53\pm 0.00$ $0.44\pm 0.00$ $0.49\pm 0.01$ $0.45\pm 0.00$ $0.42\pm 0.00$ $0.34\pm 0.00$ $0.45\pm 0.00$
PaLM1-8B $0.45\pm 0.00$ $0.51\pm 0.00$ $0.46\pm 0.00$ $0.49\pm 0.01$ $0.48\pm 0.01$ $0.44\pm 0.01$ $0.34\pm 0.00$ $0.48\pm 0.01$
PaLM1-62B $0.45\pm 0.00$ $0.50\pm 0.01$ $0.47\pm 0.00$ $0.49\pm 0.01$ $0.49\pm 0.00$ $0.44\pm 0.00$ $0.33\pm 0.00$ $0.48\pm 0.01$
PaLM1-540B $0.46\pm 0.01$ $0.49\pm 0.01$ $0.47\pm 0.00$ $0.50\pm 0.01$ $0.50\pm 0.00$ $0.44\pm 0.00$ $0.33\pm 0.01$ $0.48\pm 0.00$
PaLM2-XXS $0.44\pm 0.01$ $0.50\pm 0.00$ $0.45\pm 0.00$ $0.50\pm 0.01$ $0.48\pm 0.00$ $0.45\pm 0.00$ $0.34\pm 0.00$ $0.49\pm 0.00$
PaLM2-XS $0.45\pm 0.01$ $0.50\pm 0.01$ $0.46\pm 0.01$ $0.49\pm 0.00$ $0.48\pm 0.00$ $0.44\pm 0.00$ $0.33\pm 0.01$ $0.49\pm 0.00$
PaLM2-S $0.45\pm 0.00$ $0.49\pm 0.00$ $0.47\pm 0.00$ $0.50\pm 0.01$ $0.50\pm 0.01$ $0.44\pm 0.00$ $0.31\pm 0.00$ $0.49\pm 0.00$
PaLM2-M $0.45\pm 0.01$ $0.49\pm 0.01$ $0.48\pm 0.01$ $0.50\pm 0.01$ $0.50\pm 0.00$ $0.44\pm 0.00$ $0.29\pm 0.00$ $0.49\pm 0.01$
PaLM2-L $0.46\pm 0.01$ $0.49\pm 0.00$ $0.49\pm 0.00$ $0.50\pm 0.00$ $0.52\pm 0.00$ $0.44\pm 0.00$ $0.28\pm 0.00$ $0.49\pm 0.00$